# Asynchronous Federated Stochastic Optimization with Exact Averaging for Heterogeneous Local Objectives

### v1

@@ A.3 Main results for convex problems
$\begin{split}\|x_{s,k}-\alpha_{k}g_{i}(x_{s,k})-x^{\star}\|^{2}&=\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k}(\|d_{i}(x_{s,k})\|^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}\langle d_{i}(x_{s,k}),x^{\star}-x_{s,k}\rangle\\ &\quad-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle\\ &\leq\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle,\end{split}$
@@
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}\right]&\leq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k}-x^{\star}\|^{2}-\frac{1}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2}.\end{split}$
@@
A.4 Numerical results for iid data
@@
A.4.1 Cross-device FL
Our numerical results for $n=128$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 8 and 9, respectively. In Fig. 8 we observe that AREA and FedBuff outperform the remaining methods, with AREA being more numerically stable and yielding tighter confidence bounds. In Fig. 9, we observe that the perfomance of multiple local SGD steps speeds up significantly the convergence of AS-FedAvg. On the other hand, while AREA achieves the second fastest convergence, it also converges to a large error neighborhood of the optimal loss function value due to the large value of the stepsize.
@@ A.4.1 Cross-device FL
A.4.2 Cross-silo FL
Our numerical results for $n=4$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 10 and 11, respectively. In Fig. 10, we observe that all methods achieve comparable performance with the exception of S-FedAvg under non-uniform client update frequencies. Conversely, in Fig. 11, we observe that the execution of multiple serial SGD steps destabilizes MIFA and FedBuff, and as in the cross-device setting, boosts the convergence of AS-FedAvg.

### v2

@@ A.1.3 Main results for convex problems
$\begin{split}\|x_{s,k}-\alpha_{k}g_{i}(x_{s,k})-x^{\star}\|^{2}&=\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k}(\|d_{i}(x_{s,k})\|^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\langle d_{i}(x_{s,k}),x^{\star}-x_{s,k}\rangle-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle\\ &\leq\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle,\end{split}$
@@
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}\right]&\leq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k}-x^{\star}\|^{2}\\ &\quad-\frac{1}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2}.\end{split}$
@@
A.2 Additional numerical results for non-iid data
A.2.1 Quadratic problem
A.2.2 Cross-silo FL
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
A.3 Numerical results for iid data
@@
A.3.1 Cross-device FL
Our numerical results for $n=128$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 7 and 8, respectively. In Fig. 7 we observe that AREA and FedBuff outperform the remaining methods, with AREA being more numerically stable and yielding tighter confidence bounds. In Fig. 8, we observe that the perfomance of multiple local SGD steps speeds up significantly the convergence of AS-FedAvg. On the other hand, while AREA achieves the second fastest convergence, it also converges to a large error neighborhood of the optimal loss function value due to the large value of the stepsize.
@@ A.3.1 Cross-device FL
A.3.2 Cross-silo FL
Our numerical results for $n=4$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 9 and 10, respectively. In Fig. 9, we observe that all methods achieve comparable performance with the exception of S-FedAvg under non-uniform client update frequencies. Conversely, in Fig. 10, we observe that the execution of multiple serial SGD steps destabilizes MIFA and FedBuff, and as in the cross-device setting, boosts the convergence of AS-FedAvg.